OpenStudy (anonymous):

use y = f(x) to graph g(x) = f(x) + 1

OpenStudy (ash2326):

g(x)=f(x)+1 shift the graph of f(x) by one unit up

OpenStudy (anonymous):

how would you graph that though?

OpenStudy (anonymous):

the y = f(x)

OpenStudy (ash2326):

y=f(x) It would be given, wouldn't it?

OpenStudy (anonymous):

oh yeah it is. so what about g(x) = f(x + 1)?

OpenStudy (anonymous):

how would i graph that? thought i had to make a t-chart or something

OpenStudy (ash2326):

g(x)=f(x+1) shift the graph one step left yeah, that's the way list down the values of f(x) for each x, then find g(x)=f(x+1) values, then plot both